OpenStudy (anonymous):

Find the exact values of the six trigonometric functions of θ if θ is in standard position and the terminal side of θ is in the given quadrant and satisfies the given condition. III; bisects the quadrant

OpenStudy (anonymous):

do you mean bisects the quadrant 3?

OpenStudy (anonymous):

if so then the values of the six will be \[\sin \theta = -\sqrt{2}/2\] \[\cos \theta=-\sqrt{2}/2\] \[\tan \theta=1\] \[\cot \theta=1\] \[\csc \theta=-\sqrt2\] \[\sec \theta=-\sqrt2\]

OpenStudy (anonymous):

yep, thanks. explain pls how did you find it?

OpenStudy (anonymous):

ok first let me sketch this

OpenStudy (anonymous):

 since the angle formed in quadrant III is 90 degrees, then if we bisect that angle we'll get half of that which is 45 degrees.... so we know that in a right triangle with 45 degrees, the sides will have a ratio of 1:1:square root of 2 so let's suppose our hypotenuse is sqrt of 2 and the legs have measure of 1. Since we're in quadrant 3 both legs will have measure of -, (because at quadrant 3, x and y are both negative) 1 but the hypotenuse will still be sqrt2... so we'll have using this info, you can now compute for the 6 trig. functions of theta.. so sin theta=-1/sqrt2=-sqrt(2)/2 cos theta=-1/sqrt2=-sqrt(2)/2 tan theta=sin theta/cos theta <a trigonometric identity since sin theta=cos theta in this case, tan theta=1 cot theta=1/tan theta=1/1=1 sec theta= 1/cos theta=1/(-1/sqrt2)=-sqrt2 csc theta=1/sin theta=1/(-1/sqrt2)=-sqrt2